OpenStudy (anonymous):

find the slope of the line containing (3,-1) and (-3,1)

OpenStudy (anonymous):

The slope is determined from the following formula where "m" is the slope:\[m = \frac{ y _{1} - y _{2} }{ x _{1} - x _{2} }\]and you use your given points:\[(x _{1}, y _{1})\]and\[(x _{2}, y _{2})\]

OpenStudy (anonymous):

So, x1 = 3, y1 = -1, x2 = -3, and y2 = 1. So you just substitute those values into the given formula.

OpenStudy (anonymous):

How's this going for you? All good now?

OpenStudy (anonymous):

i got -2/0 and my choices are m=1/3 m=-1/3 m=0 and m=3 :/ did i do it wrong?

OpenStudy (anonymous):

you should get: (-1 - 1) / [3 - (-3)] = -2/(3 + 3) = -2/6 = -1/3

OpenStudy (anonymous):

Making sense now?

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

You're very welcome!

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition!

OpenStudy (anonymous):

awhh thanks, you're welcome as well! :)